Working together, Paul and Daniel can pick forty bushels of apples in 4.95 hours. Had he done it alone it would have taken Daniel 9 hours. Find how long it would take Paul to do it alone.
Answer by richwmiller(17219)

4.95/9+4.95/x=1
x=11
or
1/9+1/p=1/4.95
p=11